### v1

@@ 6.1.1 Divisive Hierarchical Clustering with "Last-mile" Training
We begin by treating the entire dataset $\hat{D}$, processed through the previous multimodal data feature representation process, as a single cluster. Using the DPC algorithm, we then partition this dataset into initial sub-clusters $\{\hat{D_{i}}\}$.
The selection of the DPC algorithm is motivated by its ability to automatically ascertain the optimal number of sub-clusters, thus eliminating the necessity for manual intervention. This capability is particularly beneficial in contexts where the exact number of clusters is not predefined. More importantly, DPC utilizes the results of LPGF where data points are more densely grouped, enhancing the clustering outcomes and ensuring more uniform data distribution within subsets.
@@
Figure 15: Calculation of $\delta$(a), Depth of Index Tree(b) and Leaf Node Characteristics for Different Dataset Sizes(c).
@@ 6.2 Index Optimization
Regarding query efficiency, different index structures are suited to different query workloads. In this section, we focus on optimizing and reordering sibling nodes (sub-clusters) that share the same parent node (parent cluster) based on the query-aware mechanism. This optimization aims to refine the index structure, thereby improving query performance and efficiency. Note that in our optimization strategy, we focus exclusively on the arrangement of all child nodes under each parent node while keeping the inheritance relationships unchanged. For instance, consider the scenario in Fig 14 where there are three non-leaf nodes ($N_{1}$,$N_{4}$,$N_{5}$) in the left branch of the cluster tree, each pointing to a list of their respective child nodes: $L_{1}=\{N_{4},N_{5}\}$, $L_{4}=\{N_{13},N_{14},N_{15}\}$, $L_{5}=\{N_{16},N_{17},N_{18}\}$. Our method is to optimize the order of nodes within each list (e.g. $L_{1}$ can be reordered as $L_{1}=\{N_{5},N_{4}\}$), without altering the parent-child relationships between nodes (e.g. $N_{5}$ cannot be moved to $L_{4}$).
@@ 7 Experiments
We implement our MQRLD index in Scala. All experiments are completed in a distributed cluster environment built on the Ubuntu Linux system, using Apache Spark and Apache Hudi as the big data computing engine and data lake management framework, respectively. The cluster consists of three nodes, each with an Intel(R) Core(TM) i7-11700F CPU and 16GB memory.
@@ 7.1.1 DataSets
OSM [^1] https://download.geofabrik.de is a spatial dataset consisting of 105 million records randomly sampled from North America in the OpenStreetMap dataset. Each record includes six attributes, such as ID, timestamp, GPS coordinates, etc. Taxi [^2] https://www1.nyc.gov/site/tlc/about/tlc- trip- record- data.page is randomly sampled from records of yellow taxi trips in New York City in 2018 and 2019. We select five attributes from the dataset: the number of passengers, trip distance, pickup, drop-off locations, and total fare, to form a five-dimensional dataset. Stock [^3] https://www.kaggle.com/datasets/ehallmar/daily-historical-stock-prices-1970-2018 contains 210 million records, consisting of daily historical stock price data for over 6,000 stocks from 1970 to 2018. We focus on four features: trading volume, opening price, high price, and adjusted closing price. Yelp [^4] https://www.yelp.com/dataset consists of 6,990,280 user reviews of business places (e.g., cafes, restaurants, hotels, shops) from 11 metropolitan areas. For each review, we extract a set of 64-dimensional semantic vectors as the experimental dataset. Color [^5] https://image-net.org/download-images is a 32-dimensional image feature dataset extracted from the ImageNet dataset, which contains 1,281,167 records. Forest [^6] https://www.kaggle.com/c/forest-cover-type-prediction/data is collected by the US Geological Survey and the US Forest Service, which includes 565,892 12-dimensional records. AI Challenger [^7] https://challenger.ai is a multimodal dataset that contains more than 300,000 images with text descriptions. In order to simulate the rich hybrid queries scenario, we define two numeric attributes for each image: the length of the text description and the size of the image file. Synthetic Datasets include datasets with three different distributions, which are Uniform, GuassMix, and Skewed. The dimensions and data volumes of these datasets are adjustable. By default, we use a 3-dimensional dataset with a volume of 10 million records.
@@ 7.1.3 Competitors
We compare our index with other high-dimensional indexing methods, as shown in Table 7. All methods store data using DataFrame structures with data buckets as the smallest physical storage unit and apply the same optimizations where applicable.
Table 7: High-dimensional Index Competitors.
@@
ZM (Wang et al., 2019) learned
ML (Davitkova et al., 2020) learned
LISA (Tian et al., 2022) learned
Qd-tree (Yang et al., 2020) learned
LIMS (Tian et al., 2022) learned
Flood (Nathan et al., 2020) learned
Tsunami (Ding et al., 2020) learned
R*-tree (Beckmann et al., 1990) tree
SPB-tree (Chen et al., 2015) tree
M-tree (Ciaccia et al., 1997) tree
@@ 7.1.4 Evaluation metrics
7.2 Evaluation of Range Query
@@
Figure 16: Range Query Evaluation Results for Different Indexes on Real Datasets.
7.3 Evaluation of KNN Query
@@
Next, we evaluate MQRLD’s performance on KNN queries using the Yelp, Color, and Forest datasets. Similarly, under the same query workload, we optimize other indexes as much as possible. It can be observed from Fig 17 that MQRLD maintains relatively low average query times across different $K$ values (due to incomparable results beyond 12 dimensions for R*-tree and LISA, we disregard them on the Yelp and Color):
@@
(1) Compared with non-learned indexes, MQRLD shows average query efficiency improvements of 4.2×, 3.1×, and 3.3× over M-tree, R*-tree, and SPB-tree respectively. This is because the leaf nodes of MQRLD contain highly similar data, and have a prediction model to improve the query speed.
@@
(2) Compared with learned indexes, MQRLD improves query efficiency by an average of 2.2×, 1.9×, and 1.5× over ML, LISA, and LIMS datasets. ML uses clustering for data mapping but suffers from prediction errors and longer search times due to its global predictive model. LISA and LIMS utilize local predictive models to improve the query efficiency for "last-mile" search, but their models are complex and the mapping process is time-consuming. In contrast, MQRLD trains simple linear regression models at each leaf node, achieving more accurate and faster queries. This is enabled by MQRLD’s data-aware and query-aware data representation and high-dimensional learned index construction strategy, which keep prediction model errors within a low threshold as well as improve MQRLD’s query efficiency.
Figure 17: KNN Query Evaluation Results for Different Indexes on Real Datasets.
7.4 Evaluation of CBR
@@
In this section, we investigate the CBR of each index on both real and synthetic datasets. Fig 18 reports the CBR for range queries across different datasets. We observe that MQRLD exhibits the lowest CBR across various datasets. The tree-based indexes, R*-tree and Qd-tree, show relatively less CBR, demonstrating that tree structure has the advantage of pruning a significant amount of irrelevant data regions. Flood uses grid indexes that also effectively filter out irrelevant data through grid partitioning, while Tsunami, which combines tree structures and grid indexes, further leverages these advantages to minimize CBR during queries. ZM and ML, however, show higher CBR across all datasets. This is due to the sequential scanning required on one-dimensional mapped values, compounded by the drawbacks of Z-order curves and prediction errors in learning models. Compared to these indexes, MQRLD combines the structural advantages of tree indexes with our index optimization strategy that ensures frequently accessed nodes are kept in positions that are more cache-friendly, allowing most queries to be completed by scanning only a few leaf nodes.
@@
Figure 18: Cross-bucket Rate Evaluation Results of Different Indexes on Different Datasets.
@@
Figure 19: Scalability Evaluation Results w.r.t. Dataset Size.
7.5 Scalability
@@
Figure 20: Scalability Evaluation Results w.r.t. Dimension.
@@
7.6 Evaluation of Rich Hybrid Queries
In this section, we evaluate the performance of rich hybrid queries. Except for MQRLD, other indexes in our experiment do not support rich hybrid queries and can only search a single vector, so we combine them to simulate the execution of rich hybrid queries. These combinations include six different hybrid indexes: M-tree $\times$ N, SPB-tree $\times$ N, ML $\times$ N, LIMS $\times$ N, LIMS + Flood, and LIMS + Tsunami. $\times$ N means a combination of N instances of the same index (N $\in$ [2,5]). Note that these combined indexes execute queries sequentially. For example, if ML $\times$ 2 is used to perform a V.K $\bigoplus$ V.R query, we will first use one ML index to execute the V.K query, then build another ML index to perform the V.R query.
@@ 7.6 Evaluation of Rich Hybrid Queries
Figure 21: Rich hybrid Query Evaluation Results for Different Indexes on Real Datasets.
@@ 7.7 Discussions
In this section, to clarify some potential concerns that have not been mentioned before, we will discuss the time and space consumption associated with the proposed MQRLD and demonstrate the effectiveness of our proposed optimization strategies through ablation experiments. Furthermore, all experiments conducted to verify these concerns utilize the GuassMix dataset as the experimental data.
@@
The Time Cost of Index Construction
For a multimodal data retrieval platform, the volume of data it stores is often substantial. This raises a concern about whether constructing the MQRLD index is highly time-consuming. Experiments demonstrate that when the dataset contains 10M data, the index construction time is 1393.6s, indicating that the index construction is not excessively time-consuming.
The Space Cost of Index
In index construction, we use a linear regression model to fit the data distribution within each leaf node, resulting in each leaf node storing only a few parameters. However, with large-scale data, the number of leaf nodes may become substantial, potentially leading to an excessively large index size despite the small number of parameters per model. Fig 22(a) demonstrates that the space complexity of the index structure we constructed on datasets with varying data volumes approximates $O(logN)$. Therefore, as the amount of multimodal data stored in the platform increases, the space cost of the MQRLD index will not grow excessively.
Figure 22: The Space Cost of MQRLD Index(a), and Ablation Study(b).
@@
To explore the impact of different components on MQRLD performance, we evaluate the performance of range queries using MQRLD in various states. Fig 22(b) provides a detailed illustration of the average query times for different query optimized methods, in which "Full Scan" refers to scanning the entire dataset to implement range query, "Initialized_MQRLD" indicates querying with the initialized MQRLD without index optimization, and "Optimized_MQRLD" represents querying with the optimized MQRLD.
@@
(1) Comparison between "Full Scan" and "Initialized_MQRLD" shows that our data representation strategy and index structure significantly reduce query times, even in the initial state.
(2) The results for "Initialized_MQRLD" and "Optimized_MQRLD" validate the effectiveness of our query-aware mechanism, with the optimized matrix $T$ and index structure further enhancing query performance.
(3) In "Optimized_MQRLD," the sampling frequency employed is the minimum effective value for both feature enhancement and index optimization, which is 10%. As the sample frequency increases, the performance of MQRLD continues to improve.

### v2

@@ 6.1.1 Divisive Hierarchical Clustering with "Last-mile" Training
We begin by treating the entire dataset $\hat{D}$, processed through the previous multimodal data feature representation process, as a single cluster. Using the DPC algorithm, we then partition this dataset into initial sub-clusters $\{\hat{D_{i}}\}$. The selection of the DPC algorithm is based on evaluations of multiple metrics, including data division time, depth of index tree, and average query time. As presented in Table 7, different clustering methods are employed in Algorithm 2. Our findings indicate that the DPC method aligns most suitably with our data division strategy. While the K-means algorithm achieves a lower data division time at K=4 compared to the DPC algorithm, it relies on a manually defined parameter K. Conversely, although the Agglomerative algorithm does not necessitate manual parameter setting, its computational complexity is higher as evidenced by significantly longer data partitioning time when compared to DPC.The DPC algorithm is capable of automatically ascertaining the optimal number of sub-clusters, and determining clusters’ centroids jointly by density and distance. Moreover, DPC utilizes the results of LPGF where data points are more densely grouped, enhancing the clustering outcomes and ensuring more uniform data distribution within subsets.
Table 7: Data division results of different cluster methods
Cluster Method Number of Clusters Data Division Time(s) Depth of Index Tree Avg Query Time(ms)
K-means K = 2 1276.92 5 72.06
K = 3 877.06 3 26.55
K = 4 611.45 2 9.11
Agglomerative - 1135.66 2 7.14
DPC - 674.54 2 6.42
  • a
    "-" indicates that the number of clusters is determined by the algorithm itself.
@@
Figure 18: Calculation of $\delta$(a), depth of index tree(b) and leaf node characteristics for different dataset sizes(c). The results from (b) and (c) demonstrate that as the dataset size increases, there is a gradual convergence in both the depth of our index tree and the number of leaf nodes, rather than a significant increase.
@@ 6.2 Index Optimization
Regarding query efficiency, different index structures are suited to different query workloads. In this section, we focus on optimizing and reordering sibling nodes (sub-clusters) that share the same parent node (parent cluster) based on the query-aware mechanism. This optimization aims to refine the index structure, thereby improving query performance and efficiency. Note that in our optimization strategy, we focus exclusively on the arrangement of all child nodes under each parent node while keeping the inheritance relationships unchanged. For instance, consider the scenario in Fig 17 where there are three non-leaf nodes ($N_{1}$,$N_{4}$,$N_{5}$) in the left branch of the cluster tree, each pointing to a list of their respective child nodes: $L_{1}=\{N_{4},N_{5}\}$, $L_{4}=\{N_{12},N_{13},N_{14}\}$, $L_{5}=\{N_{15},N_{16},N_{17}\}$. Our method is to optimize the order of nodes within each list (e.g. $L_{1}$ can be reordered as $L_{1}=\{N_{5},N_{4}\}$), without altering the parent-child relationships between nodes (e.g. $N_{5}$ cannot be moved to $L_{4}$).
@@ 7 Experiments
We implement our MQRLD index in Scala and apply it to the data lake and vector database respectively. The data lake is deployed in a distributed cluster environment consisting of three nodes, using Apache Spark as the computing engine and Apache Hudi for data lake management. Each node is equipped with 64-bit Ubuntu 18.04, Intel(R) Core(TM) i7-11700F CPU @ 2.50 GHz, and 16 GB RAM. The vector database is deployed on a single-node server using Milvus as its management framework. The server’s hardware configuration includes 64-bit Ubuntu 20.04, two Intel(R) Xeon(R) Gold 6226 CPUs @ 2.90GHz, 256 GB RAM, and an NVIDIA RTX A5000 GPU.
@@ 7.1.1 DataSets
OSM [^1] https://download.geofabrik.de is a spatial dataset consisting of 105 million records randomly sampled from North America in the OpenStreetMap dataset. Each record includes six attributes, such as ID, timestamp, GPS coordinates, etc. Taxi [^2] https://www1.nyc.gov/site/tlc/about/tlc- trip- record- data.page is randomly sampled from records of yellow taxi trips in New York City in 2018 and 2019. We select five attributes from the dataset: the number of passengers, trip distance, pickup, drop-off locations, and total fare, to form a five-dimensional dataset. Stock [^3] https://www.kaggle.com/datasets/ehallmar/daily-historical-stock-prices-1970-2018 contains 210 million records, consisting of daily historical stock price data for over 6,000 stocks from 1970 to 2018. We focus on four features: trading volume, opening price, high price, and adjusted closing price. Yelp [^4] https://www.yelp.com/dataset consists of 6,990,280 user reviews of business places (e.g., cafes, restaurants, hotels, shops) from 11 metropolitan areas. For each review, we extract a set of 64-dimensional semantic vectors as the experimental dataset. Color [^5] https://image-net.org/download-images is a 32-dimensional image feature dataset extracted from the ImageNet dataset, which contains 1,281,167 records. Forest [^6] https://www.kaggle.com/c/forest-cover-type-prediction/data is collected by the US Geological Survey and the US Forest Service, which includes 565,892 12-dimensional records. AI Challenger [^7] https://challenger.ai is a multimodal dataset that contains more than 300,000 images with text descriptions. In order to simulate the rich hybrid queries scenario, we define two numeric attributes for each image: the length of the text description and the size of the image file. SIFT1B [^8] http://corpus-texmex.irisa.fr/ is a standard benchmark dataset that contains 1 billion of 128-dimensional feature vectors. We select 0.1M-100M pieces of data from it. LAION400M [^9] https://laion.ai/ is a multimodal dataset containing 400 million image-text pairs. To generate rich hybrid queries, we concatenate two 512-dimensional feature vectors representing the image and text, along with a 2-dimensional scalar vector denoting the original dimensions of the image, resulting in 1026-dimensional vectors. Synthetic Datasets include datasets with three different distributions, which are Uniform, GuassMix, and Skewed. The dimensions and data volumes of these datasets are adjustable. By default, we use a 3-dimensional dataset with a volume of 10 million records.
@@ 7.1.3 Competitors
We compare our index with other multi-dimensional and high-dimensional indexes, as shown in Table 9. All methods store data using DataFrame structures with data buckets as the smallest physical storage unit and apply the same optimizations where applicable.
Table 9: Competitors.
@@
ZM (Wang et al., 2019) multi-dimensional
ML (Davitkova et al., 2020) multi-dimensional
LISA (Tian et al., 2022) multi-dimensional
Qd-tree (Yang et al., 2020) multi-dimensional
LIMS (Tian et al., 2022) multi-dimensional
Flood (Nathan et al., 2020) multi-dimensional
Tsunami (Ding et al., 2020) multi-dimensional
R*-tree (Beckmann et al., 1990) multi-dimensional
SPB-tree (Chen et al., 2015) multi-dimensional
M-tree (Ciaccia et al., 1997) multi-dimensional
HNSW(Malkov and Yashunin, 2018) high-dimensional
IVF(Moura and Cristo, 2009) high-dimensional
LSH(Gionis et al., 1999) high-dimensional
DB-LSH(Tian et al., 2024) high-dimensional
@@ 7.1.4 Evaluation Metrics
(2) Recall. Recall is the proportion of true relevant neighbors retrieved by a KNN query out of all the actual relevant neighbors. It measures the completeness of the search in finding the true nearest neighbors.
@@
7.2 Evaluation of Range Query for Multi-dimensional Indexes
@@
Figure 19: Range query evaluation results for multi-dimensional indexes on real datasets.
7.3 Evaluation of KNN Query for Multi-dimensional Indexes
@@
Next, we evaluate MQRLD’s performance on KNN queries using the Yelp, Color, and Forest datasets. Similarly, under the same query workload, we optimize other indexes as much as possible. It can be observed from Fig 20 that MQRLD maintains relatively low average query times across different $K$ values:
@@
(1) Compared with non-learned multi-dimensional indexes, MQRLD shows average query efficiency improvements of 4.2×, 3.1×, and 3.3× over M-tree, R*-tree, and SPB-tree respectively. This is because the leaf nodes of MQRLD contain highly similar data, and have a prediction model to improve the query speed.
@@
(2) Compared with learned multi-dimensional indexes, MQRLD improves query efficiency by an average of 2.2×, 1.9×, and 1.5× over ML, LISA, and LIMS datasets. ML uses clustering for data mapping but suffers from prediction errors and longer search times due to its global predictive model. LISA and LIMS utilize local predictive models to improve the query efficiency for "last-mile" search, but their models are complex and the mapping process is time-consuming. In contrast, MQRLD trains simple linear regression models at each leaf node, achieving more accurate and faster queries. This is enabled by MQRLD’s data-aware and query-aware data representation and high-dimensional learned index construction strategy, which keep prediction model errors within a low threshold as well as improve MQRLD’s query efficiency.
Figure 20: KNN query evaluation results for multi-dimensional indexes on real datasets. Due to incomparable results beyond 12 dimensions for R*-tree and LISA, we disregard them on the Yelp and Color.
7.4 Evaluation of CBR for Multi-dimensional Indexes
@@
In this section, we investigate the CBR of several multi-dimensional indexes on both real and synthetic datasets. Fig 21 reports the CBR for range queries across different datasets. We observe that MQRLD exhibits the lowest CBR across various datasets. The tree-based indexes, R*-tree and Qd-tree, show relatively less CBR, demonstrating that tree structure has the advantage of pruning a significant amount of irrelevant data regions. Flood uses grid indexes that also effectively filter out irrelevant data through grid partitioning, while Tsunami, which combines tree structures and grid indexes, further leverages these advantages to minimize CBR during queries. ZM and ML, however, show higher CBR across all datasets. This is due to the sequential scanning required on one-dimensional mapped values, compounded by the drawbacks of Z-order curves and prediction errors in learning models. Compared to these indexes, MQRLD combines the structural advantages of tree indexes with our index optimization strategy that ensures frequently accessed nodes are kept in positions that are more cache-friendly, allowing most queries to be completed by scanning only a few leaf nodes.
@@
Figure 21: Cross-bucket rate evaluation results of range query and KNN query for multi-dimensional indexes on different datasets. Due to incomparable results beyond 12 dimensions for R*-tree and LISA, we disregard them on Yelp and Color.
@@
Figure 22: Scalability evaluation results w.r.t. dataset size for multi-dimensional indexes.
7.5 Evaluation of Scalability for Multi-dimensional Indexes
@@
Figure 23: Scalability evaluation results w.r.t. dimension for multi-dimensional indexes. Due to incomparable results beyond 12 dimensions for R*-tree and LISA, we disregard them on experiments of 16 dimensions.
@@
7.6 Evaluation of Rich Hybrid Queries for Multi-dimensional Indexes
In this section, we evaluate the performance of rich hybrid queries for multi-dimensional indexes. Except for MQRLD, other indexes in our experiment do not support rich hybrid queries and can only search a single vector, so we combine them to simulate the execution of rich hybrid queries. These combinations include six different hybrid indexes: M-tree $\times$ N, SPB-tree $\times$ N, ML $\times$ N, LIMS $\times$ N, LIMS + Flood, and LIMS + Tsunami. $\times$ N means a combination of N instances of the same index (N $\in$ [2,5]). Note that these combined indexes execute queries sequentially. For example, if ML $\times$ 2 is used to perform a V.K $\bigoplus$ V.R query, we will first use one ML index to execute the V.K query, then build another ML index to perform the V.R query.
@@ 7.6 Evaluation of Rich Hybrid Queries for Multi-dimensional Indexes
Figure 24: Rich hybrid query evaluation results for multi-dimensional indexes.
7.7 Evaluation of KNN Query for High-dimensional Indexes
To fully evaluate the KNN query performance of MQRLD against other high-dimensional indexes on datasets of varying sizes, we select the SIFT1B dataset with sizes of 1M, 10M, 100M for comparison. We compare the average query time and the recall-time curve of different high-dimensional indexes. As shown in the left figure of Fig 25, MQRLD’s advantage in both query time and recall efficiency becomes increasingly apparent as the dataset grows larger. On the dataset of size 100M, MQRLD’s average query time is 1.3× to 2.8× faster than other high-dimensional indexes, and it takes the least time to reach the same recall. Other high-dimensional indexes’ data structures are sensitive to increasing data volumes, such as the DB-LSH and LSH hash table structures, the clustering structures included in IVF, and the HNSW single-layer graph structures, resulting in their query times increasing rapidly as the data size grows. In contrast, MQRLD’s structure is robust in the face of increasing data volume, due to the predictive models maintained in the leaf nodes that allow for rapid filtering of large amounts of data during queries. At the same time, the high-precision predictive models combined with efficient multimodal data representation strategies give MQRLD a significant advantage over other vector indexes in the time to reach the same recall on large-scale datasets.
Figure 25: KNN query evaluation results for high-dimensional indexes.
7.8 Evaluation of Rich Hybrid Queries for High-dimensional Indexes
In this section, we evaluate the performance of various high-dimensional indexes for rich hybrid queries. Except for MQRLD, other high-dimensional indexes only support vector queries. Therefore, we combine these indexes with Tsunami to evaluate rich hybrid queries, similar to Section 7.6. The experimental dataset is a subset of the LAION400M dataset with sizes of 1M, 10M, 100M. Fig 26 shows the average query time and recall-time curve for different indexes. It can be observed that MQRLD significantly outperforms other combined indexes on the 100M datasets. On average, MQRLD’s query time is 1.2× faster than the sub-optimal index DB-LSH+Tsunami, and its recall-time curve demonstrates that it consistently achieves the same recall faster than other indexes. These experimental results highlight that MQRLD not only supports rich hybrid queries but also substantially enhances query performance for high-dimensional data, especially on the largest dataset.
Figure 26: Hybrid query evaluation results for high-dimensional indexes.
@@ 7.9 Discussions
In this section, to clarify some potential concerns that have not been mentioned before, we will discuss the trade-off between query workload and data appending, time and space complexity associated with MQRLD, and demonstrate the effectiveness of our proposed optimization strategies through ablation experiments. Furthermore, all experiments conducted to verify these concerns utilize the GuassMix dataset as the experimental data.
The Trade-off Between Query Workload and Data Appending
The trade-off between multimodal data query workload (considered as read operations) and data appending (considered as write operations) is a hot topic, which is always analyzed in specific application scenarios. For multimodal data retrieval platforms, the performance of data queries has a direct impact on downstream tasks. As a result, ensuring the performance of the query workload is more critical than supporting data appending. The platform’s ability to handle query workload primarily depends on the multi-dimensional or high-dimensional indexing methods it employs, as discussed in Section 2. Most existing methods focus on achieving fast and accurate retrieval without analyzing continuous data growing (e.g., Flood, Tsunami, LISA, ML, ZM, RSMI, LMSFC, etc). Only a few studies can support the functionality of data appending (e.g., HNSW, IVFADC, E2LSH, DB-LSH, etc). Given that large-scale multimodal data is accumulated over long periods, with the data appending frequency being low and primarily occurring in batch forms, we offline established a high-dimensional learned index for MQRLD to support high-performance retrieval and preserve its functionality of data appending. Since the learned index structure is based on a cluster tree, data appending can also be achieved through search operations. By querying new data and identifying its nearest neighbor (i.e., equivalent to a KNN search, K=1), we can determine the most likely insert location for it. As shown in Fig 18 (b) and (c), the index structure changes only when the data grows exponentially. Therefore, even when data is frequently appending, the index structure remains stable. However, it is undeniable that if new excessive data accumulates, the retrieval performance will deteriorate, requiring index reconstruction when the query accuracy falls below a predefined threshold.
@@
The Time Complexity of A Single Query
Assuming a multimodal dataset contains $N$ records, the complexity of the traverse time from the root node to the leaf node is bounded by $O(logN)$, while the complexity of the "last-mile" search using the linear regression model inside the leaf node is $O(1)$. Therefore, the time complexity of a single query is bounded by $O(logN)$.
The Time Complexity of Index Construction and Experiments
In each iteration of the Divisive Hierarchical Clustering process, DPC calculates the density and distance of each data point to others in order to find cluster centers, resulting in a complexity of $O(N^{2})$. The number of iterations is bounded by $O(logN)$, so the overall time complexity for the Divisive Hierarchical Clustering phase is bounded by $O(N^{2}logN)$. Then, for each leaf node, a linear regression model is trained. The training step for all leaf nodes has a time complexity of $O(N)$. The overall time complexity of the indexing construction is dominated by the Divisive Hierarchical Clustering process, with a complexity bounded by $O(N^{2}logN)$. Fig 27(a) illustrates the index construction times for multi-dimensional and high-dimensional indexes. The multi-dimensional indexes are evaluated using the GaussMix dataset, where we observe that MQRLD’s index construction time is comparable to that of other multi-dimensional indexes. As for the high-dimensional indexes, the SIFT10M dataset is employed, and we observe that the index construction times of high-dimensional indexes are lower than multi-dimensional indexes in general.
The Space Complexity of Index and Experiments
Since our clustering algorithm ensures that intermediate nodes in the cluster tree have at least two child nodes, the cluster tree will, in the worst case, degrade into a full binary tree with $N$ leaf nodes. The total number of non-leaf nodes in such a tree is $N-1$. Leaf nodes store a tuple containing its centroid, radius, and a linear regression model, requiring $O(N)$ space. Non-leaf nodes store a tuple consisting of its centroid, radius, and a pointer, also requiring $O(N)$ space. Therefore, the total space complexity of our index is bounded by $O(N)$. However, due to the efficiency of our iterative and divisive approach, the tree depth is low, and the total number of nodes is significantly less than $12N-$ 1 (as shown in Fig 18 (b) and (c)), making our index structure practically very small. Fig 27(b) shows the index size of the index structure we constructed on datasets with varying data volumes. We can see that MQRLD saves an average of 65.7%-90.1% space compared to multi-dimensional indexes and an average of 81.4%-99.7% compared to high-dimensional indexes.
Figure 27: The time cost of different index construction(a), The space cost of different index construction(b), and ablation study(c).
Experiments of Deploying MQRLD on Vector Database
In this section, we apply MQRLD to the vector database Milvus and compare its KNN query performance with high-dimensional indexes HNSW, IVF, LSH, and DB-LSH. The experimental dataset uses the SIFT1B dataset with sizes of 0.1M, 1M, 10M. Due to the limitations of the Milvus database in our experiment environment, we cannot use larger datasets to evaluate query performance. The left figure of Fig 28 shows the average query time for MQRLD and different high-dimensional indexes at various dataset sizes. We can see that as the data scale increases, the advantage of MQRLD gradually becomes apparent. On the dataset of size 10M, MQRLD’s average query time is significantly lower than IVF and LSH. The increase in data scales of the same order of magnitude does not significantly impact the tree depth of MQRLD’s cluster tree, thereby resulting in a negligible increase in MQRLD’s query time. The right figure of Fig 28 shows the recall-time curves for MQRLD and different high-dimensional indexes at various dataset sizes. Compared to other indexes, MQRLD takes the least time to achieve a 100% recall.
Figure 28: Evaluation results for high-dimensional indexes on vector database.
@@
To explore the impact of different components on MQRLD performance, we evaluate the performance of range queries using MQRLD in various states. Fig 27(b) provides a detailed illustration of the average query times for different query optimized methods, in which "Full Scan" refers to scanning the entire dataset to implement range query, "Initialized_MQRLD" indicates querying with the initialized MQRLD without index optimization, "Optimized_T" indicates querying with MQRLD optimized after feature representation, and "Optimized_Index" indicates querying with MQRLD optimized using index structures. It is noteworthy that the sampling frequency used at different optimization stages is the minimum effective value, which is 10%. We believe that as the sampling frequency increases, the performance of MQRLD will continue to improve.
@@
(1) The results of "Full Scan" and "Initialized_MQRLD" shows that our data representation strategy and index structure significantly reduce query times, even in the initial state.
(2) The results of "Initialized_MQRLD" and "Optimized_T" validate the effectiveness of our feature representation method, with the optimized matrix T further enhancing query performance.
(3) The results of "Optimized_T" and "Optimized_Index" validate the effectiveness of our index structure, with the optimized index structure effectively reducing the scanning of irrelevant data buckets and reducing query time.
(4) Overall experimental results demonstrate the effectiveness of our query-aware mechanism, with the optimized matrix T and index structure further enhancing query performance.